OpenStudy (anonymous):

What is the acceleration of a ball traveling horizontally with an initial velocity of 20 meters/second and, 2.0 seconds later, a velocity of 30 meters/second?

OpenStudy (fifciol):

Do you know the relation between acceleration and velocity?

OpenStudy (fifciol):

\[a=\frac{ \Delta v }{ \Delta t }=\frac{ v-v_0 }{ t-t_0 }\]

OpenStudy (fifciol):

\[v-v_0 \] is change in velocity \[t-t_0 \] is change in time

OpenStudy (fifciol):

so v-v0 =30 -20 =10 t-t0 = 2-0=2 what is a then?